Part of a circuit is shown. If the voltage across
Answer #1

given

resistors R, 10 ohm are connected in parallel combination
and the voltage across 10 ohm resistor is 3 V, current through the circuit is 0.9 A

as the resistors are connected in parallel so the current through resistors will be divided


now current through 10 ohm resistor from ohm's law I = V/R = 3/10 = 0.3 A

so remaining 0.6 A current will be through resistor R,and the potential difference across R, 10 ohm reistor is same

   I= V/R => 0.6 = 3/R ==> R = 3/0.6 = 5 ohm

the resistance of the resistor is 5 ohm